# Efficient designs for multivariate crossover trials

### v1

@@ 4 Illustration using the gene data
We go back to use the genetic data example from Leaker et al., (2017) to evaluate design efficiency under different covariance structures. In particular, we focus on the $3\times 3$ crossover design used in the gene data example, involving $18$ subjects and $3$ treatment sequences, namely $ABC$, $CAB$ and $BCA$. Subjects $1-6$ receive treatment sequence $ABC$, $7-12$ receive $CAB$, while remaining receive $BCA$, and multiple gene expression values are measured from each subject in the three time periods. Note the study design is a uniform design with $p=t=3$. We name this design $d_{0}$, where $d_{0}$ belongs to $\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$. Next, we evaluate the performance of $d_{0}$ under various correlation structures and compare it with other possible designs.
@@ 4.1 Proportional structure
Under the proportional structure, to investigate trace optimality/efficiency of $d_{0}$ we consider two different structures of matrix $\bm{V}$ as follows:
@@
    $\bm{V}=\left(1-r_{(1)}^{2}\right)^{-1}\begin{bmatrix}1&r_{(1)}&r_{(1)}^{2}\\ r_{(1)}&1&r_{(1)}\\ r_{(1)}^{2}&r_{(1)}&1\end{bmatrix}$, where $-1<r_{(1)}<1$;
@@
    $\bm{V}=\begin{bmatrix}1&r_{(1)}&0\\ r_{(1)}&1&r_{(1)}\\ 0&r_{(1)}&1\end{bmatrix}$, where $-1/\sqrt{2}<r_{(1)}<1/\sqrt{2}$.
Structure 1 corresponds to a $AR(1)$ structure, while structure 2 corresponds to a tridiagonal structure. The constraint on the values of $r_{(1)}$ is to ensure positive definiteness of the matrix $\bm{V}$. From Theorem 3.2 in Subsection 3.1, we know that under the proportional structure, a design $d^{*}\in\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$ represented by $OA_{I}\left(n=\lambda t\left(t-1\right),p=t,t,2\right)$, where $\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$ is a class of binary designs with $p=t$, $\lambda$ is a positive integer and $t\geq 3$, is trace optimal/efficient for the direct effects over $\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$ for the multivariate response case. For any binary design $d\in\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$, with $t$ and $n$ same as that of $d^{*}$, let $e$ represent the efficiency of design $d$, defined as (Kunert, 1991)
@@
Since $d^{*}$ maximizes $tr\left(\bm{C}_{d(s1)}\right)$, over $\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$, $e$ lies between 00 and $1$, and values of $e$ close to $1$ indicate that $d$ is a highly efficient design.
@@
$e$ $=\frac{tr\left(\bm{\varGamma}^{-1}\right)tr\left(\bm{C}_{d(uni)}\right)}{tr\left(\bm{\varGamma}^{-1}\right)tr\left(\bm{C}_{d^{*}(uni)}\right)}=\frac{tr\left(\bm{C}_{d(uni)}\right)}{tr\left(\bm{C}_{d^{*}(uni)}\right)},$ (4.1)
@@
[…]
(a) $\bm{V}$: $AR(1)$ Structure
(b) $\bm{V}$: Tridiagonal Structure
Figure 4: Efficiency of $d_{0}$.
@@ 4.2 Generalized markov-type structure
To check the trace optimality of design $d_{0}$ and compare its performance with $d^{*}$ under the Markovian structure, we consider $4$ different cases based on the structure of matrices $\bm{V}_{1}$ and $\bm{V}_{R}$. These cases are same as considered in Subsubsection 3.2.1. Note that here $d^{*}$ is a design represented by an orthogonal array of Type $I$ and strength $2$ with $p=t=3$ and $n=18$. Here, we also assume that $\sigma_{11}=\sigma_{22}$. Figure 5 displays $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d_{0}}$, $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{*}}$, $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d_{0}}$ and $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{*}}$ under the $4$ cases, corresponding to various non-zero permissible values of $r_{(1)}$. Here, $RD_{d_{0}}$ and $RD_{d^{*}}$ is as given in (3.16) with $d^{(0)}=d_{0},d^{*}$. Note that lower values of $RD_{d^{(0)}}$ indicate that $d^{(0)}$ is a highly efficient design. From Figure 5 and for all cases, $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d_{0}}$ and $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{*}}$ never attains 00. So, $d_{0}$ and $d^{*}$ fails to attain the upper bound $u$. Also, $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d_{0}}$ is quite higher than $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{*}}$ for all cases, so we may say that $d^{*}$ performs better than $d_{0}$. Taking all the $4$ cases in account, the maximum value of $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d_{0}}$ and $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{*}}$ is around $99.02\%$ and $4\%$, respectively, indicating the very poor performance of $d_{0}$ when compared to $d^{*}$. From these comparisons we may conclude that if the experiment for the gene study had used an orthogonal array based design then he/she would have scored much in terms of efficiency. The poor choice of designs may also effect the parameter estimation for the gene dataset.
(a) $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{0},d^{*}\}$ (Case $1$)
(b) $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{0},d^{*}\}$ (Case $1$)
(c) $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{0},d^{*}\}$ (Case $2$)
(d) $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{0},d^{*}\}$ (Case $2$)
(e) $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{0},d^{*}\}$ (Case $3$)
(f) $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{0},d^{*}\}$ (Case $3$)
(g) $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{0},d^{*}\}$ (Case $4$)
(h) $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{0},d^{*}\}$ (Case $4$)
Figure 5: Plotting $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ and $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{0},d^{*}\}$, where $p=t=3$, $n=18$ and $\sigma_{11}=\sigma_{22}$.
5 Conclusion and future direction
@@
In this article, we studied the trace optimality/efficiency of an orthogonal array of Type $I$ and strength $2$ when measurements are recorded on $g$ responses,where $g\geq 1$. To account for correlation between responses, we considered two different types of structure for the error dispersion matrix, namely, the proportional structure and the generalized markov-type structure. The main conclusions of this paper are:
@@
    Under the generalized markov-type structure, for the $4$ different cases depending on the choices of matrices $\bm{V}_{1}$ and $\bm{V}_{R}$, when $p=t=3$ and $n=6$, and $p=t=4$ and $n=12$, an orthogonal array of type $I$ and strength $2$ is highly efficient for the direct effects over the class of binary designs.
Although in this article, we have only considered universal optimality and trace optimality criteria, other criteria can also be investigated. In the future, for multivariate crossover designs, we plan to investigate $A$-, $D$-, and $E$-optimal design, under the proportional and markovian structures of the dispersion matrix.
Statements and Declarations
Funding
The authors declare that no financial support has been received from any funding agency, for this work.
Competing Interests
The authors declare that there are no conflicts of interest regarding the publication of this article.
Ethical Conduct
This research does not involve human or animal participants.
Authors’ Contributions
All authors contributed to the study conception and design. Material preparation, analysis and investigation were performed by Shubham Niphadkar. The first draft of the manuscript was written by Shubham Niphadkar and all authors commented on previous versions of the manuscript. All authors read and approved the final manuscript.
@@ A.1  Proportional structure
$\bm{C}_{d(s1)}=\bm{\varGamma}^{-1}\otimes\\ \left[\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{V}^{*}\right)\bm{T}_{d}-\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{V}^{*}\right)\bm{F}_{d}\left(\bm{F}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{V}^{*}\right)\bm{F}_{d}\right)^{-}\bm{F}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{V}^{*}\right)\bm{T}_{d}\right],$ (A.2)
where $\bm{\varGamma}$ is a $g\times g$ positive definite and symmetric matrix with non-zero off-diagonal elements. Thus from (A.1) and (A.2), we can prove (3.3). ∎
@@
$\bm{C}_{d^{*}(s1)}$ $=\bm{\varGamma}^{-1}\otimes\bm{C}_{d^{*}(uni)},$ (A.3)
where $\bm{C}_{d^{*}(uni)}=\bm{C}_{d^{*}(uni)(11)}-\bm{C}_{d^{*}(uni)(12)}\bm{C}_{d^{*}(uni)(22)}^{-}\bm{C}_{d^{*}(uni)(21)}$, and $\bm{\varGamma}$ is a $g\times g$ positive definite and symmetric matrix with non-zero off-diagonal elements. Here, $\bm{C}_{d^{*}(uni)(11)}=\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{V}^{*}\right)\bm{T}_{d^{*}}$, $\bm{C}_{d^{*}(uni)(12)}=\bm{C}_{d^{*}(uni)(21)}^{{}^{\prime}}=\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{V}^{*}\right)\bm{F}_{d^{*}}$, $\bm{C}_{d^{*}(uni)(22)}=\bm{F}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{V}^{*}\right)\bm{F}_{d^{*}}$ and the matrix $\bm{V}^{*}=\bm{V}^{-1}-\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}^{-1}\bm{{1}}_{p}\right)^{-1}\bm{V}^{-1}\bm{J}_{p\times p}\bm{V}^{-1}$.
@@
$tr\left(\bm{T}^{{}^{\prime}}_{d^{*}1}\bm{\psi}^{{}^{\prime}}\bm{V}^{*}\bm{\psi}\bm{T}_{d^{*}1}\right)-\frac{\left(\bm{V}^{*}\right)_{1,1}}{t}$ and $|\bm{E}|\neq 0$. Here, $\left(\bm{V}^{*}\right)_{1,1}$ is the element corresponding to first row and first column of the matrix $\bm{V}^{*}$, and $\bm{T}_{d^{*}1}$ is the treatment incidence matrix for the first subject corresponding to $d^{*}$. For $g>1$, let us denote
$\bm{\varGamma}^{-1}=\begin{bmatrix}\gamma^{(11)}&\ldots&\gamma^{(1g)}\\ \vdots&\vdots&\vdots\\ \gamma^{(g1)}&\ldots&\gamma^{(gg)}\end{bmatrix}.$
Since $\bm{\varGamma}$ is a $g\times g$ positive definite, symmetric and non-diagonal matrix, the same holds true for $\bm{\varGamma}^{-1}$. So, from (A.3) and (A.4), we get that $\bm{C}_{d^{*}(s1)}$ is completely symmetric for $g>1$ case if and only if
@@
where $\gamma^{(kk^{\prime})}$ is the $(k,k^{\prime})^{th}$ element of the matrix $\bm{\varGamma}^{-1}$. Here, due to positive definiteness, symmetricity and non-diagonality of $\bm{\varGamma}^{-1}$, $\gamma^{(kk)}>0$ and for atleast one pair $(k,k^{\prime})$, $\gamma^{(kk^{\prime})}=\gamma^{(k^{\prime}k)}\neq 0$. Thus (A.5) does not hold true. Hence, $\bm{C}_{d^{*}(s1)}$ is not a completely symmetric matrix for $g>1$ case. ∎

### v2

@@ 4 Illustration using the gene data
We go back to use the genetic data example from Leaker et al., (2017) and evaluate the design used under different covariance structures. In particular, we focus on the $3\times 3$ crossover design used in the gene data example, involving $18$ subjects and $3$ treatment sequences, namely $ABC$, $CAB$ and $BCA$. Subjects $1$-$6$ receive treatment sequence $ABC$, $7$-$12$ receive $CAB$, while the remaining receive $BCA$ and multiple gene expression values are measured from each subject in three time periods. Note the study design is a uniform design with $p=t=3$. We name this design $d_{0}$, where $d_{0}$ belongs to $\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$, with $t=3$ and $\lambda=3$.
@@ 4.1 Proportional structure
Under the proportional structure, to investigate trace optimality/efficiency of $d_{0}$ we consider three different structures of matrix $\bm{V}$ as discussed in Table 1. Corresponding to each case, for $p=3$, the matrix $\bm{V}$ is as follows:
@@
    $\bm{V}=\begin{bmatrix}1&r_{(1)}&r_{(1)}^{2}\\ r_{(1)}&1&r_{(1)}\\ r_{(1)}^{2}&r_{(1)}&1\end{bmatrix}$, where $10<r_{(1)}<$ 1;
@@
    $\bm{V}=\begin{bmatrix}1&\left[1+log\left(r_{(1)}\right)\right]r_{(1)}&\left[1+2log\left(r_{(1)}\right)\right]r_{(1)}^{2}\\ \left[1+log\left(r_{(1)}\right)\right]r_{(1)}&1&\left[1+log\left(r_{(1)}\right)\right]r_{(1)}\\ \left[1+2log\left(r_{(1)}\right)\right]r_{(1)}^{2}&\left[1+log\left(r_{(1)}\right)\right]r_{(1)}&1\end{bmatrix}$, where $10<r_{(1)}<$ 1;
  3. 3.
    $\bm{V}=\begin{bmatrix}1&r_{(1)}&r_{(1)}^{4}\\ r_{(1)}&1&r_{(1)}\\ r_{(1)}^{4}&r_{(1)}&1\end{bmatrix}$, where $10<r_{(1)}<$ 1.
From Theorem 3.2 in Subsection 3.1, we know that under the proportional structure, a design $d^{*}\in\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$ represented by $OA_{I}\left(n=\lambda t\left(t-1\right),p=t,t,2\right)$, where $\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$ is a class of binary designs with $p=t$, $\lambda$ is a positive integer and $t\geq 3$, is trace optimal/efficient for the direct effects over $\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$ for the multivariate response case. For any binary design $d\in\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$, with $t$ and $n$ same as that of $d^{*}$, let $e$ represent the efficiency of design $d$, defined as (Kunert, 1991)
@@
Since $d^{*}$ maximizes $tr\left(\bm{C}_{d(s1)}\right)$ over $\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$, $e$ lies between 00 and $1$, and values of $e$ close to $1$ indicate that $d$ is a highly efficient design.
@@
$e$ $=\frac{tr\left(\bm{\mathit{\Gamma}}^{-1}\right)tr\left(\bm{C}_{d(uni)}\right)}{tr\left(\bm{\mathit{\Gamma}}^{-1}\right)tr\left(\bm{C}_{d^{*}(uni)}\right)}=\frac{tr\left(\bm{C}_{d(uni)}\right)}{tr\left(\bm{C}_{d^{*}(uni)}\right)},$ (4.1)
@@
From (4.1), it is clear that for a particular design $d\in\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$, and fixed values of $t$, $n$ and $p$, $e$ depends only on the matrix $\bm{V}$, thus only on $r_{(1)}$. Our computations for $10<r_{(1)}<$ 1, show that the maximum value of $e$ under the three structures of $\bm{V}$ is around $2.78\%$, indicating that $d_{0}$ performs very poorly in comparison to the efficient design $d^{*}$.
@@ 4.2 Generalized Markov-type structure
To check the trace optimality of design $d_{0}$ and compare its performance with $d^{*}$ under the Markovian structure, we consider cases $5$, $6$ and $7$ from Subsubsection 3.2.1. Note that here $d^{*}$ is a design represented by an orthogonal array of Type $I$ and strength $2$ with $p=t=3$ and $n=18$. We assume that $\sigma_{11}=\sigma_{22}$. Fig. 8 displays $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d_{0}}$, $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{*}}$, $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d_{0}}$ and $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{*}}$ under the $3$ cases considered. From Fig. 8, we note $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d_{0}}$ and $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{*}}$ never attain 00, and $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d_{0}}$ and $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{*}}$ are around $99.70\%$ and $0.38\%$, respectively, indicating the poor performance of $d_{0}$ when compared to $d^{*}$. From these comparisons, we may conclude that if the experiment for the gene study had used an orthogonal array based design then he/she would have scored much in terms of efficiency. The poor choice of designs may also affect the parameter estimation for the gene dataset.
(a) $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{0},d^{*}\}$
($\bm{V}_{1}$: Mat($\infty$), $\bm{V}_{R}$: Mat($0.5$))
(b) $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{0},d^{*}\}$
($\bm{V}_{1}$: Mat($\infty$), $\bm{V}_{R}$: Mat($0.5$))
(c) $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{0},d^{*}\}$
($\bm{V}_{1}$: Mat($\infty$), $\bm{V}_{R}$: Mat($1.5$))
(d) $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{0},d^{*}\}$
($\bm{V}_{1}$: Mat($\infty$), $\bm{V}_{R}$: Mat($1.5$))
(e) $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{0},d^{*}\}$
($\bm{V}_{1}$: Mat($0.5$), $\bm{V}_{R}$: Mat($0.5$))
(f) $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{0},d^{*}\}$
($\bm{V}_{1}$: Mat($0.5$), $\bm{V}_{R}$: Mat($0.5$))
Figure 8: Plots of $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ and $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{0},d^{*}\}$, where $p=t=3$, $n=18$ and $\sigma_{11}=\sigma_{22}$.
5 Computational details
The exploratory analysis for the gene data regarding the presence of varying effects, and within and between response correlation (see Section 1.1), was illustrated through line plots and correlation plots with the p-values of the correlation test, respectively. Also, the comparison of designs (see Subsubsection 3.2.1 and Section 4) is represented using line plots. For the above purpose, we used R programming (R Core Team, 2023) with R version 4.3.1 under Windows 11 (64-bit), with a 13th Gen Intel(R) Core(TM) I5-1335U processor and 8GB of RAM. We used base, MASS, pracma, raster, rapportools, writexl, graphics, dplyr, tidyr, ggplot2, readxl, snpar, stats, QuantPsyc, corrplot packages. The R codes (.R files), the dataset on the $3$ genes (.zip file), and the MATLAB code (.m file) used for cross-checking in Subsubsection 3.2.1 are available at https://github.com/rsphd/Efficient-designs-for-multivariate-crossover-trials.
6 Conclusion
@@
In this article, we studied the trace optimality/efficiency of an orthogonal array of Type $I$ and strength $2$ when measurements are recorded on $g$ responses, where $g>1$. To account for correlation between responses, we considered two different types of structure for the error dispersion matrix, namely, the proportional structure and the generalized Markov-type structure. Under the proportional covariance structure, we considered within response covariance structures, i.e, structures of $\bm{V}_{1}$ to be Mat($0.5$), Mat($1.5$) and Mat($\infty$). Under the generalized Markov-type covariance structure, we considered $7$ different choices for structure of ($\bm{V}_{1}$, $\bm{V}_{R}$), namely (Mat($0.5$), Mat($1.5$)), (Mat($0.5$), Mat($\infty$)), (Mat($1.5$), Mat($0.5$)), (Mat($1.5$), Mat($\infty$)), (Mat($\infty$), Mat($0.5$)) and (Mat($\infty$), Mat($1.5$)). The main conclusions of this paper are:
@@
    Under the generalized Markov-type structure, for the $7$ different cases depending on the choices of matrices $\bm{V}_{1}$ and $\bm{V}_{R}$, when $p=t=3$ and $n=6$, and $p=t=4$ and $n=12$, an orthogonal array of type $I$ and strength $2$ is highly efficient for the direct effects over the class of binary designs.
Although in this article, we have only considered universal optimality and trace optimality criteria, other criteria can also be investigated. In the future, for multivariate crossover designs, we plan to investigate $A$-, $D$-, and $E$-optimal designs, under the proportional and Markovian structures of the dispersion matrix.
@@ A.1  Proportional structure
$\bm{C}_{d(s1)}=\bm{\mathit{\Gamma}}^{-1}\otimes\\ \left[\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{V}^{*}\right)\bm{T}_{d}-\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{V}^{*}\right)\bm{F}_{d}\left(\bm{F}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{V}^{*}\right)\bm{F}_{d}\right)^{-}\bm{F}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{V}^{*}\right)\bm{T}_{d}\right],$ (A.2)
where $\bm{\mathit{\Gamma}}$ is a $g\times g$ positive definite and symmetric matrix with non-zero off-diagonal elements. Thus from (A.1) and (A.2), we can prove (3.3). ∎
@@
$\bm{C}_{d^{*}(s1)}$ $=\bm{\mathit{\Gamma}}^{-1}\otimes\bm{C}_{d^{*}(uni)},$ (A.3)
where $\bm{C}_{d^{*}(uni)}=\bm{C}_{d^{*}(uni)(11)}-\bm{C}_{d^{*}(uni)(12)}\bm{C}_{d^{*}(uni)(22)}^{-}\bm{C}_{d^{*}(uni)(21)}$, and $\bm{\mathit{\Gamma}}$ is a $g\times g$ positive definite and symmetric matrix with non-zero off-diagonal elements. Here, $\bm{C}_{d^{*}(uni)(11)}=\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{V}^{*}\right)\bm{T}_{d^{*}}$, $\bm{C}_{d^{*}(uni)(12)}=\bm{C}_{d^{*}(uni)(21)}^{{}^{\prime}}=\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{V}^{*}\right)\bm{F}_{d^{*}}$, $\bm{C}_{d^{*}(uni)(22)}=$
$\bm{F}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{V}^{*}\right)\bm{F}_{d^{*}}$ and the matrix $\bm{V}^{*}=\bm{V}^{-1}-\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}^{-1}\bm{{1}}_{p}\right)^{-1}\bm{V}^{-1}\bm{J}_{p\times p}\bm{V}^{-1}$.
@@
$tr\left(\bm{T}^{{}^{\prime}}_{d^{*}1}\bm{\psi}^{{}^{\prime}}\bm{V}^{*}\bm{\psi}\bm{T}_{d^{*}1}\right)-\frac{\left(\bm{V}^{*}\right)_{1,1}}{t}$ and $|\bm{E}|\neq 0$. Here $\left(\bm{V}^{*}\right)_{1,1}$ is the element corresponding to the first row and first column of the matrix $\bm{V}^{*}$, and $\bm{T}_{d^{*}1}$ is the treatment incidence matrix for the first subject corresponding to $d^{*}$. For $g>1$, let us denote
$\bm{\mathit{\Gamma}}^{-1}=\begin{bmatrix}\gamma^{(11)}&\cdots&\gamma^{(1g)}\\ \vdots&\vdots&\vdots\\ \gamma^{(g1)}&\cdots&\gamma^{(gg)}\end{bmatrix}.$
Since $\bm{\mathit{\Gamma}}$ is a $g\times g$ positive definite, symmetric and non-diagonal matrix, the same holds for $\bm{\mathit{\Gamma}}^{-1}$. So from (A.3) and (A.4), we get that $\bm{C}_{d^{*}(s1)}$ is completely symmetric for $g>1$ case if and only if
@@
where $\gamma^{(kk^{\prime})}$ is the $(k,k^{\prime})^{th}$ element of the matrix $\bm{\mathit{\Gamma}}^{-1}$. Here, due to positive definiteness, symmetricity and non-diagonality of $\bm{\mathit{\Gamma}}^{-1}$, $\gamma^{(kk)}>0$ and for atleast one pair $(k,k^{\prime})$, $\gamma^{(kk^{\prime})}=\gamma^{(k^{\prime}k)}\neq 0$. Thus (A.5) does not hold. Hence, $\bm{C}_{d^{*}(s1)}$ is not a completely symmetric matrix for $g>1$ case. ∎